OpenStudy (anonymous):

Use the appropriate rules to determine teh derivitaive of.... i'll write teh equation below w/ teh equation tool

OpenStudy (anonymous):

\[f(x) = 3x ^{5} - 2x ^{3/2} - (7/x ^{5}) + \sqrt{93}\]

OpenStudy (anonymous):

well first thing to think of is that you can take the derivative of each term separately and then add them together.

OpenStudy (anonymous):

so what's the derivative of 3x^5 to start?

OpenStudy (anonymous):

well i have to do it first

OpenStudy (anonymous):

isnt there a simpler way to do it with teh rules?

OpenStudy (anonymous):

I'm using the rules: the derivate of a sum is the sum of the derivatives of the terms.

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

anyway, you can use the rules to differentiate 3x^5

OpenStudy (anonymous):

you subtract one from the power and multiply by the power: 5*3x^4

OpenStudy (anonymous):

can u show me the steps?

OpenStudy (anonymous):

to differentiate 3x^5 you multiply by the power of x, 5, and then subtract one from the power of x, so it becomes 5-1=4, so in total we have 5*3x^4

OpenStudy (anonymous):

\[(ax^n)'=nax^{n-1}\]

OpenStudy (anonymous):

a=3, n=5 in this case.

OpenStudy (anonymous):

so its 3(5x^4) - 2(x^1/2) -..how about teh rest??

OpenStudy (anonymous):

the derivative of the second term is -3x^(1/2)

OpenStudy (anonymous):

-2*(3/2)=-3

OpenStudy (anonymous):

so there are still two terms left to do.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

im not sure how to approach teh other 2

OpenStudy (anonymous):

-(7/x^5)=-7x^-5

OpenStudy (anonymous):

now you can do it just like before.

OpenStudy (anonymous):

so taht will turn into -7(-5x^-6) or 35x^-6)

OpenStudy (anonymous):

right

OpenStudy (anonymous):

do i just leave teh sqrt of 93 or what do i do with it?

OpenStudy (anonymous):

sqrt(93) is a constant, because there's no x in it, the derivative of a constant is zero.

OpenStudy (anonymous):

oh ok so so far i have: f'(x) = 15x^4 - 3x^1/2 + 35x^-6

OpenStudy (anonymous):

is taht the answer

OpenStudy (anonymous):

that's right

OpenStudy (anonymous):

tahnk you soo much!! can u help me with limits??

OpenStudy (anonymous):

not now, sorry, I'm going to bed.

OpenStudy (anonymous):

oh ok but thanks 4 ur help

OpenStudy (anonymous):

You're welcome